OpenStudy (kingswank):

Use the net to find the surface area of the cylinder.Diameter= 4yd and dHeight= 12y

OpenStudy (gorv):

A=2*pi*r^2+2*pi*r*h

OpenStudy (gorv):

is it 4yd???

OpenStudy (kingswank):

@gorv yeah

OpenStudy (gorv):

okay wait

OpenStudy (kingswank):

@gorv ok

OpenStudy (gorv):

diameter = 2*r radius=diameter /2 surface area=2*pi*(2y)^2+2*pi*2yd*12y =25.13y^2*y^2+150.7y^2 =175.92y^2

OpenStudy (kingswank):

@gorv Thanks again bro

OpenStudy (gorv):

welcome